OpenStudy (anonymous):

6x-y=-6 determine the x- and y-intercepts

OpenStudy (anonymous):

The x-intercept is where the line crosses the x-axis, or where y=0. Therefore, 6x-0=-6 (sub in zero for y) x=-1 (divide by the coefficient) So the x-intercept is (-1,0) The y-intercept is almost identical, but you instead have x=0 6(0)-y=-6 (sub in x=0) y=6 (divide by coefficient) and you receive the y-intercept (0,-6)